# One Model, Any Conjunctive Query: Graph Neural Networks for Answering Queries over Incomplete Knowledge Graphs

### v2

@@ 6.4 Ablation studies
How does AnyCQ perform outside the training domain? As mentioned in Section˜5.4, we expect the search engine to exhibit similar behavior on processed instances, regardless of the underlying knowledge graph. We validate this claim by applying AnyCQ models trained on FB15k-237 or on NELL to both datasets, equipping a relevant link predictor. The results on our QAR benchmarks are presented in Table˜4. Notably, the differences between models’ accuracies in QAR are marginal, confirming that the resulting search engine is versatile and not dataset-dependent. In fact, the model trained on FB15k-237 exhibits better performance on both datasets, further aligning with our assumption on the transferability and generalizability of AnyCQ.
How does AnyCQ perform with a prefect link predictor? The AnyCQ framework’s performance heavily depends on the underlying link prediction model, responsible for guiding the search and determining the satisfiability of generated assignments. Hence, to assess purely the quality of our search engines, we equipped them with perfect link predictors for the test KGs, eliminating the impact of predictors’ imperfections. The results of experiments on our QAR and QAC benchmarks are available in Table˜4 and Table˜5 , respectively. Remarkably, the simple query types in QAC pose no challenge for AnyCQ, which achieves $100\%$ F1-score on all of them. Furthermore, on large formula splits, the F1-score remains over $90\%$, displaying the accuracy of our search framework. Similarly, for the task of QAR, AnyCQ with a perfect link predictor achieved over $90\%$ F1-score, establishing the engine’s excellent ability to retrieve answers to structurally complex questions.
How does AnyCQ’s scale when the query size increases? Let $Q=\exists\vec{y}.\Phi(\vec{y})$ be a conjunctive Boolean query over a knowledge graph $G$. Denote by $|Q|$ the number of literals in $Q$, let $h$ be the maximum arity of a literal in $Q$, and let $\vec{c}=(c_{0},...,c_{s})$ be the mentioned constants. Then, the corresponding computational graph contains $(|V(G)|+1)\cdot|\vec{y}|+2\cdot|\vec{c}|+|Q|$ vertices and at most $|V(G)|\cdot(|\vec{y}|+h\cdot|Q|)+h\cdot|\vec{c}|\cdot|Q|$ edges. Since AnyCQ processes this graph in linear time, the computational complexity of a single search step is:
Importantly, this complexity does not depend on the structure of the query graph of $Q$ and scales only linearly with the sizes of the input formula and the KG $G$. We validate this linearity, evaluating the average AnyCQ search step time for queries of different sizes from the ‘3hub’ splits. The results, presented in Figure˜5, indicate that the empirical performance matches the theoretical analysis. In particular, Figure˜5 b) and c) show that the processing time, divided by the number of variables $|\vec{y}|$ or the number of literals $|Q|$ in the input query, respectively, does not grow as the size of the query increases. We even notice a slight decreasing trend, which we attribute to efficient GPU accelerations. The difference between step times on FB15k-237-QAR and NELL-QAR remains consistent with the relative sizes of the underlying knowledge graphs. A further analysis of processing time, including a comparison with the SQL engine, is provided in Appendix˜E.
@@ 7 Summary, limitations, and outlook
In this work, we devise and study two new tasks from the query answering domain: query answer classification and query answer retrieval. Our formulations target the challenge of classifying and generating answers to structurally complex formulas with an arbitrary number of free variables. Moreover, we introduce datasets consisting of instances beyond the processing capabilities of existing approaches, creating strong benchmarks for years to come. To address this demanding setting, we introduce AnyCQ, a framework applicable for scoring and generating answers for large conjunctive formulas with arbitrary arity over incomplete knowledge graphs. We demonstrate the effectiveness over our QAC and QAR benchmarks, showing that on simple samples, AnyCQ matches the performance of state-of-the-art CQA models, while setting challenging baselines for the large instance splits. One potential limitation is considering by default the input query in disjunctive normal form, converting to which may require exponentially many operations. We hope our work will motivate the field of query answering to recognize the classification nature of the induced tasks and expand the scope of CQA to previously intractable cases.
Appendix A AnyCQ details
A.1 Architecture
@@
AnyCQ’s architecture is based on the original ANYCSP Tönshoff et al. [2023] framework. The trainable components of the AnyCQ GNN model $\theta$ are:
@@
  • a GRU Chung et al. [2014] cell $\mathbf{G}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with a trainable initial state $\mathbf{h}\in\mathbb{R}^{d}$
@@ A.2 Training methodology
For the training data, we use the training splits of the existing FB15k-237 and NELL CQA datasets Ren and Leskovec [2020], consisting of queries of types: ‘1p’, ‘2p’, ‘3p’, ‘2i’, ‘3i’, ‘2in’, ‘3in’, ‘pin’, ‘inp’ (see Table˜6 for the corresponding first-order logic formulas). Hence, during training, AnyCQ witnesses queries with projections, intersections and negations, learning principles of this logical structures. However, all of these queries mention at most 3 free variables, remaining limited in size.
@@
A.3 Hyperparameters and implementation
Architecture. We choose the hidden embedding size $d=128$ in the AnyCQ architecture for all experiments. All MLPs used in our model consist of two fully connected layers with ReLU Agarap [2018] activation function. The intermediate dimension of the hidden layer is chosen to be 128.
Training. The REINFORCE Williams [2004] discount factor $\lambda$ is set to $0.75$ for both datasets, following the best configurations in ANYCSP experiments. During training, we run our models for $T_{\text{train}}=15$ steps. The batch size is set to $4$ for FB15k-237 and $1$ for NELL, due to the GPU memory constraints. All models are trained with an Adam Kingma and Ba [2015] optimizer with learning rate $5\cdot 10^{-6}$ on a single NVIDIA Tesla V100 SXM2 with 32GB VRAM. We let the training run for 4 days, which translates to 500,000 batches on FB15k-237 and 200,000 batches for NELL, and choose the final model for testing.
@@ A.3 Hyperparameters and implementation
A.4 Scope of formulas
@@
A.5 Expressivity
Standard graph neural networks are known to have limited expressive power Xu et al. [2019], e.g. MPNNs cannot produce different outputs for graphs not distinguishable by the Weisfeiler-Lehman algorithm Leman and Weisfeiler [1968]. We argue that AnyCQ does not suffer from this limitation. It has been noticed that including randomness in GNN models increases their expressiveness Abboud et al. [2021]. In our case, for any Boolean conjunctive query $Q=\exists\vec{y}\Phi(\vec{y})$ over a knowledge graph $G$ and a relevant link predictor $\pi$, for any assignment $\alpha:\{\vec{y}\}\to V(G)$, there is a non-zero probability of $\alpha$ being selected at some point of the search (see Section˜A.1). Hence, any AnyCQ model has a chance of correctly predicting $S_{\pi,G}(Q)$, making it fully expressive for the tasks of QAC and QAR.
A.6 Fuzzy logic
Fuzzy logic has been widely adopted in the CQA literature as a way to evaluate the satisfiability of logical formulas in a continuous, differentiable manner. It underpins several prominent methods, including CQD [Arakelyan et al., 2020], GNN-QE [Zhu et al., 2022], and QTO [Bai et al., 2023], due to its modularity and interpretability. However, especially when applied to large and structurally complex queries, fuzzy logic introduces several limitations that should be taken into account.
@@ A.6 Fuzzy logic
Gradient instability in supervised learning. As discussed by Van Krieken et al. [2022], another issue with fuzzy logic arises in differentiable learning settings, where gradients must propagate through the query structure and the fuzzy connectives. This can lead to vanishing or unstable gradients, especially for large or cyclic queries. In our case, however, this problem is largely avoided: AnyCQ is trained using reinforcement learning, where the fuzzy logic score is used as a scalar reward signal and not differentiated through. During training, we apply REINFORCE, which treats the Boolean score as an external reward, and during inference, fuzzy logic is only used to rank complete assignments. As such, our framework sidesteps the gradient-related challenges described in Van Krieken et al. [2022], while retaining the benefits of fuzzy logic for scoring.
A.7 Edge labels
@@
A.7.1 LE Labels: Guiding Local Improvements
Light edge (LE) labels were originally introduced in the ANYCSP [Tönshoff et al., 2023] framework as the primary mechanism for guiding discrete search. In the context of query answering, their purpose is to identify marginal changes to the current variable assignment that are likely to increase the number of satisfied literals in the query. That is, given an assignment $\alpha$, LE labels help determine which single-variable substitutions are most promising for improving the current solution. This makes them particularly useful during local exploitation, where the goal is to refine an existing candidate assignment rather than explore the full space.
@@ A.7.1 LE Labels: Guiding Local Improvements
A.7.2 PE Labels: Prioritizing Promising Assignments
@@
and the literal $\psi_{2}=s(y_{1},y_{2})$. Let $a_{2}\in\mathcal{D}(y_{2})$ be a viable assignment to $y_{2}$. To evaluate the PE label $P_{E}(v_{\psi_{2}},v_{y_{2}\rightarrow a_{2}})$, we check whether $\exists y_{1}.s(y_{1},a_{2})$ is satisfied, i.e. whether there exists some $a_{1}$ such that the literal $s(a_{1},a_{2})$ is true, according to the link predictor $\pi$. If such $a_{1}$ exists, we set the label to 1, and otherwise - to 0. This allows the GNN to prioritize reasoning about value assignments that could plausibly contribute to satisfying the query, rather than wasting capacity on highly unlikely candidates.
@@
Our main alternative bases on the closed world assumption (CWA) Libkin and Sirangelo [2009], which restricts the set of entities that should be considered for prediction of unobserved facts. Formally, let $G$ be an observable knowledge graph and let $\tilde{G}$ be its completion. Then, for any $r\in R(G)$ and any $a,b\in V(G)$:
@@
An alternative to relying on an additional set of labels to prevent the search from accessing unreasonable assignments could be restricting the domains $\mathcal{D}(y)$ of the considered variables. In the current formulation, each variable $y$ mentioned in the input Boolean query $Q$ is assigned a domain $\mathcal{D}(y)=V(G)$. Reducing the considered domains can significantly shrink the computational graph, leading to faster computation. Such a solution would be specifically beneficial when operating on large knowledge graphs, and even essential for applications to milion-scale KGs, such as Wikidata-5M [Wang et al., 2021]. While this approach improves inference efficiency, improper application can render correct answers unreachable due to excessively restrictive domain reductions. Consequently, we leave further exploration of this direction for future work.
Appendix B Proofs
Theorem 5.1.
@@ Appendix B Proofs
Proposition B.1 (Scores of a Perfect Link Predictor).
@@
By Proposition˜B.1, this implies:
@@
Appendix C Dataset constructions
Benchmark datasets in the existing query answering literature, FB15k-237 Toutanova et al. [2015] and NELL Carlson et al. [2010], comprise formulas with simple structures, thereby impeding the comprehensive evaluation and advancement of methodologies and algorithms. We address this gap by creating new validation and test datasets on top of well-established benchmarks, consisting of queries with complexity exceeding the processing power of known approaches. In particular, we increase the number of variables mentioned in the considered formulas from 3 to between 12 and 20, while imposing structural difficulty by sampling query graphs with multiple cycles, long-distance reasoning steps and multi-way conjunctions.
C.1 Base large query generation

### v3

@@ 6.3 Ablation studies
How does AnyCQ perform outside the training domain?
As mentioned in Section˜5.4, we expect the search engine to exhibit similar behavior on processed instances, regardless of the underlying knowledge graph. We validate this claim by applying AnyCQ models trained on FB15k-237 or on NELL to both datasets, equipping a relevant link predictor. The results on our QAR and QAC benchmarks are presented in Tables˜3 and 4. respectively. Notably, the differences between models’ accuracies in QAR are marginal, confirming that the resulting search engine is versatile and not dataset-dependent. In fact, the model trained on FB15k-237 exhibits better performance on both datasets, further aligning with our assumption on the transferability and generalizability of AnyCQ.
How does AnyCQ perform with a prefect link predictor?
The AnyCQ framework’s performance heavily depends on the underlying link prediction model, responsible for guiding the search and determining the satisfiability of generated assignments. Hence, to assess purely the quality of our search engines, we equipped them with perfect link predictors for the test KGs, eliminating the impact of predictors’ imperfections. The results of experiments on our QAR and QAC benchmarks are available in Table˜3 and Table˜4, respectively. The performance of AnyCQ on all QAR splits exceeds 90%, displaying great accuracy of the GNN-guided search engine. We witness similar results on the large formula splits in QAC datasets, further confirming the relevance of our model and highlighting the impact of the equipped link predictor. Remarkably, the simple query types in QAC pose no challenge for AnyCQ, which achieves $100\%$ F1-score on all of them, with a single exception.
@@ 7 Summary, limitations, and outlook
In this work, we devise and study two new tasks from the query answering domain: query answer classification and query answer retrieval. Our formulations target the challenge of classifying and generating answers to structurally complex formulas with an arbitrary number of free variables. Moreover, we introduce datasets consisting of instances beyond the processing capabilities of existing approaches, creating strong benchmarks for years to come. To address this demanding setting, we introduce AnyCQ, a framework applicable for scoring and generating answers for large conjunctive formulas with arbitrary arity over incomplete knowledge graphs. We demonstrate the effectiveness over our QAC and QAR benchmarks, showing that on simple samples, AnyCQ matches the performance of state-of-the-art CQA models, while setting challenging baselines for the large instance splits. One potential limitation is considering by default the input query in disjunctive normal form, converting to which may require exponentially many operations. We hope our work will motivate the field of query answering to expand the scope of CQA to previously intractable cases and recognize the classification nature of the induced tasks.
Acknowledgements
The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. https://doi.org/10.5281/zenodo.22558
Appendix A Extended discussion of task formulations and baselines
A.1 Limitations of existing problem formulations
Intractability of high-arity query evaluation. The objective of complex query answering is to rank all possible answers to a given logical formula. Already for queries $Q(x_{1},x_{2})$ with two free variables, this entails scoring $|V(G)|^{2}$ pairs of entities $(a_{1},a_{2})\in V(G)^{2}$, which is computationally infeasible for modern knowledge graphs [fb15k237, nell] containing thousands of nodes. As a result, most of the existing approaches are not designed to handle higher arity queries, either resolving to inefficient enumeration strategies [fit] or approximating answers by marginal predictions. This scalability bottleneck has already been observed by [efok_dataset], who suggested more tractable evaluation methodologies, yet again being only marginal approximations of the true performance. Therefore, we argue that the ranking-based formulation has significantly limited the progress in query answering over formulas with multiple free variables.
Limited structural complexity in existing benchmarks. A related limitation lies in the structural simplicity of existing benchmarks [betae]. Standard CQA literature predominantly focuses on tree-like queries, which aligns with the capabilities of most current models. More recently, efok_dataset introduced a dataset containing cyclic queries and queries with up to two free variables; however, the overall structures remained constrained – featuring at most four variables and a single cycle. We argue that addressing structurally richer queries is essential for advancing automated reasoning systems. In real-world applications to autonomous systems, such as an AI trip planner that simultaneously books flights, accommodations, and activities while satisfying budget and availability constraints, the underlying reasoning involves large, complex queries with multiple variables. As AI agents become more capable, the complexity of the queries they must resolve is only expected to increase, requiring more expressive answering engines, e.g. applicable to large, cyclic, multi-variate formulas.
Lack of probabilistic calibration in ranking-based methods. Practical applications often demand binary decisions - answering questions like “Is X true?” or “What is the correct answer to Y?”, requiring models to classify candidate solutions as either true or false [cqa_classification]. However, many ranking-based CQA methods do not natively support this decision-making paradigm, as they focus on ordering candidates without enforcing a meaningful threshold to distinguish valid answers from incorrect ones. While many of these models are trained using classification losses, such training does not guarantee the output scores correspond to calibrated satisfiability probabilities. In fact, several approaches rely on Noisy Contrastive Estimation [lmpnn, clmpt] or apply ad hoc score-to-probability transformations [qto, fit], further weakening the reliability of predicted scores in downstream tasks.
A.2 Limitations of existing baselines
As mentioned in Section˜6.1, for the large-query splits of both QAC and QAR, none of the existing approaches are directly applicable. In particular, no method can simultaneously (1) efficiently handle cyclic queries and (2) provide calibrated probability estimates without relying on knowledge of trivial answers.
Standard methods like BetaE [betae], CQD [cqd], ConE [cone], GNN-QE [gnn-qe] or QTO [qto] are limited to tree-like queries. Neural approaches, such as LMPNN [lmpnn] or CLMPT [clmpt], are trained using Noisy Contrastive Estimation; hence, their predictions do not meaningfully translate to desired probabilities. Finally, FIT [fit] and Q2T [q2t] require transforming scores predicted by their ComplEx-based link predictors, while all known schemes (see Section˜E.1) assume the set of easy answers is known, or otherwise, trivial to recover.
Appendix B AnyCQ details
B.1 Architecture
@@
AnyCQ’s architecture is based on the original ANYCSP [anycsp] framework. The trainable components of the AnyCQ GNN model $\theta$ are:
@@
  • a GRU [gru] cell $\mathbf{G}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with a trainable initial state $\mathbf{h}\in\mathbb{R}^{d}$
@@ B.2 Training methodology
For the training data, we use the training splits of the existing FB15k-237 and NELL CQA datasets [betae], consisting of queries of types: ‘1p’, ‘2p’, ‘3p’, ‘2i’, ‘3i’, ‘2in’, ‘3in’, ‘pin’, ‘inp’ (see Table˜5 for the corresponding first-order logic formulas). Hence, during training, AnyCQ witnesses queries with projections, intersections and negations, learning principles of this logical structures. However, all of these queries mention at most 3 free variables, remaining limited in size.
@@
B.3 Hyperparameters and implementation
Architecture. We choose the hidden embedding size $d=128$ in the AnyCQ architecture for all experiments. All MLPs used in our model consist of two fully connected layers with ReLU [agarap2018deep] activation function. The intermediate dimension of the hidden layer is chosen to be 128.
Training. The REINFORCE [reinforce] discount factor $\lambda$ is set to $0.75$ for both datasets, following the best configurations in ANYCSP experiments. During training, we run our models for $T_{\text{train}}=15$ steps. The batch size is set to $4$ for FB15k-237 and $1$ for NELL, due to the GPU memory constraints. All models are trained with an Adam [adam] optimizer with learning rate $5\cdot 10^{-6}$ on a single NVIDIA Tesla V100 SXM2 with 32GB VRAM. We let the training run for 4 days, which translates to 500,000 batches on FB15k-237 and 200,000 batches for NELL, and choose the final model for testing.
@@ B.3 Hyperparameters and implementation
B.4 Trained AnyCQ  versus random search
To clarify the contribution of the training procedure in AnyCQ, we conducted an additional ablation comparing the fully trained model against an untrained random-search baseline. In this case, each variable assignment is chosen independently and uniformly at random during the search process. The resulting F1 scores, evaluated on QAC benchmarks, are presented in Table 6. Unsurprisingly, the performance drops drastically when training is removed. In particular, the random search fails completely on large-query splits, where the model must correctly assign values to over eight variables.
That said, we would like to acknowledge the potential of “simpler” strategies, such as greedy search or hill-climbing, as promising future directions. While such algorithms were not applicable in the ranking-based formulation of prior work (which required full entity rankings), our classification-based setup now makes them feasible. Exploring these strategies could yield competitive baselines for complex query answering; however, we consider this beyond the scope of the current study.
Table 6: F1-scores of trained and randomly initialzed models on QAC datasets.
Dataset Model 2p 3p pi ip inp pin 3-hub 4-hub 5-hub
FB15k-237-QAC random 3.3 0.0 4.9 2.9 4.4 3.5 0.0 0.0 0.0
AnyCQ 75.8 71.3 82.1 78.8 76.7 75.7 52.4 49.9 51.9
NELL-QAC random 2.8 0.0 3.3 2.4 3.1 3.1 0.0 0.0 0.0
AnyCQ 76.2 72.3 79.0 75.4 76.7 75.3 57.2 52.6 58.2
B.5 Scope of formulas
@@
B.6 Expressivity
Standard graph neural networks are known to have limited expressive power [xu18], e.g. MPNNs cannot produce different outputs for graphs not distinguishable by the Weisfeiler-Lehman algorithm [leman1968reduction]. We argue that AnyCQ does not suffer from this limitation. It has been noticed that including randomness in GNN models increases their expressiveness [gnn-rni]. In our case, for any Boolean conjunctive query $Q=\exists\vec{y}\Phi(\vec{y})$ over a knowledge graph $G$ and a relevant link predictor $\pi$, for any assignment $\alpha:\{\vec{y}\}\to V(G)$, there is a non-zero probability of $\alpha$ being selected at some point of the search (see Section˜B.1). Hence, any AnyCQ model has a chance of correctly predicting $S_{\pi,G}(Q)$, making it fully expressive for the tasks of QAC and QAR.
B.7 Fuzzy logic
Fuzzy logic has been widely adopted in the CQA literature as a way to evaluate the satisfiability of logical formulas in a continuous, differentiable manner. It underpins several prominent methods, including CQD [cqd], GNN-QE [gnn-qe], and QTO [qto], due to its modularity and interpretability. However, especially when applied to large and structurally complex queries, fuzzy logic introduces several limitations that should be taken into account.
@@ B.7 Fuzzy logic
Gradient instability in supervised learning. As discussed by van2022analyzing, another issue with fuzzy logic arises in differentiable learning settings, where gradients must propagate through the query structure and the fuzzy connectives. This can lead to vanishing or unstable gradients, especially for large or cyclic queries. In our case, however, this problem is largely avoided: AnyCQ is trained using reinforcement learning, where the fuzzy logic score is used as a scalar reward signal and not differentiated through. During training, we apply REINFORCE, which treats the Boolean score as an external reward, and during inference, fuzzy logic is only used to rank complete assignments. As such, our framework sidesteps the gradient-related challenges described in van2022analyzing, while retaining the benefits of fuzzy logic for scoring.
B.8 Edge labels
@@
B.8.1 LE Labels: Guiding Local Improvements
Light edge (LE) labels were originally introduced in the ANYCSP [anycsp] framework as the primary mechanism for guiding discrete search. In the context of query answering, their purpose is to identify marginal changes to the current variable assignment that are likely to increase the number of satisfied literals in the query. That is, given an assignment $\alpha$, LE labels help determine which single-variable substitutions are most promising for improving the current solution. This makes them particularly useful during local exploitation, where the goal is to refine an existing candidate assignment rather than explore the full space.
@@ B.8.1 LE Labels: Guiding Local Improvements
B.8.2 PE Labels: Prioritizing Promising Assignments
@@
and the literal $\psi_{2}=s(y_{1},y_{2})$. Let $a_{2}\in\mathcal{D}(y_{2})$ be a viable assignment to $y_{2}$. To evaluate the PE label $P_{E}(v_{\psi_{2}},v_{y_{2}\rightarrow a_{2}})$, we check whether $\exists y_{1}.s(y_{1},a_{2})$ is satisfied, i.e. whether there exists some $a_{1}$ such that the literal $s(a_{1},a_{2})$ is true, according to the link predictor $\pi$. If such $a_{1}$ exists, we set the label to 1, and otherwise – to 0. This allows the GNN to prioritize reasoning about value assignments that could plausibly contribute to satisfying the query, rather than wasting capacity on highly unlikely candidates.
@@
Our main alternative bases on the closed world assumption (CWA) [open_world_assumption], which restricts the set of entities that should be considered for prediction of unobserved facts. Formally, let $G$ be an observable knowledge graph and let $\tilde{G}$ be its completion. Then, for any $r\in R(G)$ and any $a,b\in V(G)$:
@@
An alternative to relying on an additional set of labels to prevent the search from accessing unreasonable assignments could be restricting the domains $\mathcal{D}(y)$ of the considered variables. In the current formulation, each variable $y$ mentioned in the input Boolean query $Q$ is assigned a domain $\mathcal{D}(y)=V(G)$. Reducing the considered domains can significantly shrink the computational graph, leading to faster computation. Such a solution would be specifically beneficial when operating on large knowledge graphs, and even essential for applications to milion-scale KGs, such as Wikidata-5M [wikidata5m]. While this approach improves inference efficiency, improper application can render correct answers unreachable due to excessively restrictive domain reductions. Consequently, we leave further exploration of this direction for future work.
Appendix C Theorems and proofs
Theorem C.1.
@@ Appendix C Theorems and proofs
Proposition C.2 (Scores of a Perfect Link Predictor).
@@
By Proposition˜C.2, this implies:
@@
Appendix D Dataset constructions
Benchmark datasets in the existing query answering literature, FB15k-237 [fb15k237] and NELL [nell], comprise formulas with simple structures, thereby impeding the comprehensive evaluation and advancement of methodologies and algorithms. We address this gap by creating new validation and test datasets on top of well-established benchmarks, consisting of queries with complexity exceeding the processing power of known approaches. In particular, we increase the number of variables mentioned in the considered formulas from 3 to between 12 and 20, while imposing structural difficulty by sampling query graphs with multiple cycles, long-distance reasoning steps and multi-way conjunctions.
D.1 Base large query generation